OpenStudy (amorfide):

Differentiate step by step I don't remember how to do it (4lnx-3)/(4lnx+3)

OpenStudy (amorfide):

would i do (vdu-udv)/v²

OpenStudy (amorfide):

@cwrw238

OpenStudy (amorfide):

@Hero can you help?

zepdrix (zepdrix):

Are you allowed to simplify your answer before you differentiate? :) Because it's currently written as a pair of conjugates, if we multiply them together it'll simplify very nicely.

OpenStudy (amorfide):

no do it straight off use substitute v and u if needed

OpenStudy (amorfide):

actually i do not think method matters xD

zepdrix (zepdrix):

Ok :) In that case, it looks like we have the product of two functions involving x. So we'll need to apply the product rule. The one that you posted above is the quotient rule :O we don't want to you that formula. \[\huge (uv)'=u'v+uv'\] This is the one we want :)

OpenStudy (amorfide):

i am sure it is quotient,...

zepdrix (zepdrix):

Oh i didn't see the division symbol in there sorry, it was hidden for some reason lolol

OpenStudy (amorfide):

i have the answer no method answer is 24/x(4lnx+3)²

zepdrix (zepdrix):

I hate that tiny text :)

zepdrix (zepdrix):

k you have the right formula then, just need help setting it up? :o

OpenStudy (amorfide):

well i know which is v and u i dont know the differential of 4lnx -3

zepdrix (zepdrix):

Recall this:\[(\ln x)'=\frac{ 1 }{ x }\] So let's apply that to our u. \[\large u=4\ln x - 3\]\[\large u'=4\frac{ 1 }{ x } - 0\] The 4 is just a constant, so we can ignore that while we differentiate.

OpenStudy (amorfide):

I feel dumb, thank you <3

zepdrix (zepdrix):

heh :3

OpenStudy (amorfide):

well.. i get 

OpenStudy (amorfide):

if i multiply by x/x will that give me the right answer?

OpenStudy (amorfide):

@zepdrix

OpenStudy (amorfide):

so how would i find the gradient of (4lnx-3)/(4lnx+3)

zepdrix (zepdrix):

 Hmmm I'm confused how you get 24/x on the top :O